OpenStudy (anonymous):

just check..3. The formula to find the surface area of a sphere is to take the radius of the sphere and square it and multiply that amount by pi and then multiply that by 4. This is written algebraically as S=4πr2, where S is the surface area, r is the radius, and the value of π is 3.14. Rewrite the equation to solve for the radius of the sphere if you know the sphere’s surface area. Then estimate the radius of a sphere given a surface area of 500 meters2. answer 4 pi r^2 = 500 4 * 3.14 r^2 = 500 12.56 r^2 = 500 r^2 = 500/12.56 = 39.8 take the squareroot, then r = sqrt(39.8) =

OpenStudy (anonymous):

hahaha ok but i got this of the internet from 2 different web sites

OpenStudy (mertsj):

You didn't solve the equation for r as the directions specified.

OpenStudy (anonymous):

can u help me out

OpenStudy (mertsj):

\[S=4\pi r^2\] 1. Divide both sides by 4 pi 2. Take the square root of both sides.

OpenStudy (anonymous):

can you write out the whole quation for me so i can solve

OpenStudy (anonymous):

im clue less

OpenStudy (mertsj):

Do you see the equation I posted?

OpenStudy (anonymous):

is that it to the equation?

OpenStudy (mertsj):

Write this on your paper:

OpenStudy (mertsj):

\[S=4\pi r^2\]

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Now draw a line under the S and another line under the 4pir^2

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Under both lines put 4pi

OpenStudy (mertsj):

Now it looks like this:

OpenStudy (mertsj):

\[\frac{S}{4\pi}=\frac{4\pi r^2}{4\pi}\]

OpenStudy (mertsj):

Cancel the 4pi on the right since 4pi divided by 4pi is 1

OpenStudy (mertsj):

So now it looks like this: \[\frac{S}{4\pi}=r^2\]

OpenStudy (mertsj):

Now take the square root of both sides.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did that

OpenStudy (mertsj):

Now replace S with 500 and find r

OpenStudy (anonymous):

500/12.56= 39.80 is that right